OpenStudy (anonymous):

Solve the initial value problem (dx/dt)-4x=cos (3t) with x(0)=0

OpenStudy (jamesj):

There are a couple of ways to skin this cat. Remember first that for the equation without the initial condition, the general solution is x(t) = x_h(t) + x_p(t) where x_h is the solution of the homogeneous equation and x_p is to the inhomogeneous equation above. Hence there are three steps 1. Find the homogeneous solution 2. Find the particular solution 3. Write down the general solution and then apply the initial condition Can you do these three steps?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i cant do the 1 step

OpenStudy (anonymous):

i need help with that part

OpenStudy (jamesj):

So the homogeneous equation is x' - 4x = 0 Is it not clear how to solve that?

OpenStudy (jamesj):

i.e., dx/dt = 4x

OpenStudy (anonymous):

but is not = to 0

OpenStudy (anonymous):

there is cos 3t

OpenStudy (jamesj):

Like I said above, step 1 is the homogeneous equation where it IS equal to zero. Step 2 is the inhomogeneous equation where it is equal to cos(3t)

OpenStudy (anonymous):

ok then ?

OpenStudy (jamesj):

So can you solve the homogeneous equation? What is that solution, x_h?

OpenStudy (anonymous):

can u expaline more coz i miss that class so i have no idea ho to solve this kind of problem

OpenStudy (jamesj):

here, watch this good lecture on the topic. This will get you up to speed and yes, it's worth the 50 minutes: http://ocw.mit.edu/courses/mathematics/18-03-differential-equations-spring-2010/video-lectures/lecture-7-first-order-linear-with-constant-coefficients/

OpenStudy (jamesj):

actually ... only 41! A bargain.

OpenStudy (anonymous):

ok thinks but can u solve this problem with me coz i wana make sure if i get the right unsewer

OpenStudy (jamesj):

You should learn the theory first. You might want to go back to lecture 3 first: http://ocw.mit.edu/courses/mathematics/18-03-differential-equations-spring-2010/video-lectures/lecture-3-solving-first-order-linear-odes/

OpenStudy (jamesj):

Especially if you can't solve the equation dx/dt - 3x = 0 This is the most basic sort of ODE there is, and is dealt with in the lecture 3 I just linked to.

OpenStudy (anonymous):

ok im kind of getting the theory and i ll try to apply on this problem but i told u i need to make sure if my unser is right or no

OpenStudy (jamesj):

IF you can't solve dx/dt - 4x = 0, then you're not going to understand my solution to your general equation and more importantly, you will completely fail your examination.

OpenStudy (anonymous):

i get for (dx/dt)-3x=0 e^tx(dx/dt)-3e^(-3t)y =0

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

?

OpenStudy (anonymous):

r u here ?

OpenStudy (jamesj):

So what exactly is the function x_h for dx/dt - 4x = 0? (I know I've changed the 3 to 4)

OpenStudy (jamesj):

I'll show you if dx/dt = 4x then \[\frac{dx}{x} = 4 \ dt\]

OpenStudy (anonymous):

ok

OpenStudy (jamesj):

Now integrate both sides to get ln x = 4t + c Finally, take the exp of both sides to get x(t) = exp(4t + c) = A.exp(4t) where A = e^c, an arbitrary constant I.e., the general solution of the homogeneous equation x' - 4 x = 0 is x(t) = Ae^(4t)

OpenStudy (jamesj):

Now for the inhomogeneous part of the solution which you need, you really, really should watch lecture 3.

OpenStudy (anonymous):

ok

OpenStudy (jamesj):

Let me know when you have the integrating factor for your equation.